OpenStudy (anonymous):

integrate ((0.25+(x^2))^0.5)

OpenStudy (solomonzelman):

substitute u for (0.25+x^2)

OpenStudy (anonymous):

but i try it . it will be more complex

OpenStudy (loser66):

we have formula for this, why don't you apply? \[\int{\sqrt{u^2\pm a^2}}du = \dfrac{u}{2}\sqrt{u^2\pm a^2}\pm\dfrac{a^2}{2}ln|u+\sqrt{u^2\pm a^2}|\] you case is + only, and u ^2 = 0.5^2 = 0.25 Dat sit

OpenStudy (anonymous):

thank you